(2x-7)(x^2-9)(2x+5)=91. Anybody plzz answer this question

Grade:10

1 Answers

safe india by sonu yadav
6 years ago
vaibhav gupta
you are a 10 class student and this not in you syllabus but one way to solve this is bylong multiplication method which can be used.
first multiply the equation to get a four degree equation which is   4x4-4x3-71x2+36x+224=0   ------(i)
on long division of (i) by its solution  (x-4) we get 4x3+12x2-23x-56=0    --------(ii)  and 4 became its root.
now divide (ii) by  (2x+7)   we get   2x2-x-8 =0  ----(iii)  and another root is  -7/2.
now solve (iii) and we get twoo roots x= (1+√65)/4,  x=(1-√65)/4.
so its roots are
  1.  x = 4
  2.  x = -7/2 = -3.500
  3.  x =(1-√65)/4=-1.766
  4.  x =(1+√65)/4= 2.266 .